OpenStudy (anonymous):

2x^4/10y^2*5y^3/4x^3 how do i multiply and state the restrictions on the variables?

OpenStudy (anonymous):

4x^7/25y^5

OpenStudy (anonymous):

the answer is xy/4 i was hoping someone knew how to get that

OpenStudy (anonymous):

Was I wrong?

OpenStudy (mertsj):

\[\frac{2x^4}{10y^2}\times\frac{5y^3}{4x^3}=\frac{10x^4y^3}{40y^2x^3}=\frac{xy}{4}\]

OpenStudy (anonymous):

how do ypu know to leave the 4 on bottom?

OpenStudy (mertsj):

because: \[\frac{10}{40}=\frac{1}{4}\]

OpenStudy (mertsj):

Also I forgot to say that x and y cannot be 0 because then the denominator would be

OpenStudy (mertsj):

nd the fractions would be undefined.

OpenStudy (anonymous):

thank you for explaining!

OpenStudy (mertsj):

yw